# Towards Reliable Forgetting: A Survey on Machine Unlearning Verification, Challenges, and Future Directions

### v1

@@ Abstract.
With growing demands for privacy protection, security, and legal compliance (e.g., GDPR), machine unlearning has emerged as a critical technique for ensuring the controllability and regulatory alignment of machine learning models. However, a fundamental challenge in this field lies in effectively verifying whether unlearning operations have been successfully and thoroughly executed. Despite a growing body of work on unlearning techniques, verification methodologies remain comparatively underexplored and often fragmented. Existing approaches lack a unified taxonomy and a systematic framework for evaluation. To bridge this gap, this paper presents the first structured survey of machine unlearning verification methods. We propose a taxonomy that organizes current techniques into two principal categories—behavioral verification and parametric verification—based on the type of evidence used to assess unlearning fidelity. We examine representative methods within each category, analyze their underlying assumptions, strengths, and limitations, and identify potential vulnerabilities in practical deployment. In closing, we articulate a set of open problems in current verification research, aiming to provide a foundation for developing more robust, efficient, and theoretically grounded unlearning verification mechanisms.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
As global concerns over data privacy continue to grow, particularly with the enforcement of regulations such as the General Data Protection Regulation (GDPR) (reg, 2018), managing data throughout its lifecycle has become a pressing compliance challenge. In response, machine unlearning has emerged as a key technology for protecting data sovereignty. By selectively removing specific data from trained models, it allows individuals to meaningfully exercise their right to be forgotten by enabling models to forget the influence of that data. For example, in facial recognition systems, when a user invokes their “right to be forgotten,” traditional deletion methods merely remove stored records. In contrast, machine unlearning goes further by erasing the data’s influence from the model itself. Similarly, in intelligent customer service systems, when a user requests the deletion of chat history, the unlearning mechanism updates the model to forget the relevant interactions, ensuring that future responses are no longer shaped by the removed content.
@@
Existing machine unlearning techniques can be broadly classified into exact unlearning (Bourtoule et al., 2021; Yan et al., 2022; Kumar et al., 2022; Dukler et al., 2023; Golatkar et al., 2023; Chowdhury et al., 2024) and approximate unlearning (Felps et al., 2020; Chen et al., 2023; Tarun et al., 2023; Gou et al., 2021; Phuong and Lampert, 2019). Exact unlearning typically relies on retraining the model to fully eliminate the influence of specific data, which is computationally expensive. To improve the practicality of unlearning in real-world applications, a range of approximate methods have been proposed. These approaches strike a balance between effectiveness and efficiency, enabling models to approximate forgetting at a significantly lower computational cost.
@@
While algorithmic methods for machine unlearning offer mechanisms to erase specific training data from models, concerns about trust and transparency persist—particularly due to the opaque, server-side nature of most machine learning services. A central and unresolved question remains: how can model providers convincingly demonstrate that the targeted data has been thoroughly and irreversibly removed?
@@
In response, a diverse range of verification techniques has emerged in recent years, drawing from disciplines such as watermarking and fingerprinting (Zheng et al., 2022; Chai and Chen, 2022; Guo et al., 2020; An et al., 2024), privacy auditing (Tu et al., 2024; Di et al., 2024; Sharma et al., 2024), and performance-based diagnostics (Graves et al., 2021; Mehta et al., 2022; Chundawat et al., 2023a, b; Kurmanji et al., 2023). These approaches aim to evaluate whether a model retains residual information after an unlearning operation. However, the absence of a standardized verification framework has led to troubling inconsistencies. For instance, Jeon et al. (2024a) demonstrate that a model may be deemed “successfully unlearned” under one verification criterion while simultaneously failing another—revealing a disconnect that undermines confidence in current practices.
@@
Overview. This paper presents a systematic overview of existing verification techniques for machine unlearning. The structure of this paper is illustrated in Figure 1. Specifically, Section 2 introduces the formal foundations of unlearning verification, including the definitions of exact and approximate unlearning, the construction of verification predicates, and the roles and interactions involved in the verification workflow. In Section 3, we review empirical methods that assess forgetting based on observable model behavior, while Section 4 focuses on formal techniques that rely on internal model information and offer theoretical guarantees. Section 5 presents a comparative analysis of these approaches, examining their respective strengths, limitations, and applicability. Section 6 discusses key reliability challenges in real-world verification scenarios, including training stochasticity, distribution shifts, and adversarial circumvention. Finally, Section 7 outlines open problems in the field and highlights promising directions for future research.
@@
(Zhao et al., 2023) Overview of FU
(Shaik et al., 2024) Overview of MU
(Romandini et al., 2024) Overview of FU
(Xu, 2024) Overview of MU and LLM-U
(Li et al., 2024a) Overview of MU
(Zhang et al., 2023) Overview of MU
(Said et al., 2023) Overview of GU
(Geng et al., 2025) Overview of LLM-U
(Chen et al., 2025) Privacy and Security in MU
(Mercuri et al., 2022) Overview of MU
(Liu et al., 2024) Overview of GenAI-U
(Wang et al., 2024c) Overview of MU
(Jaman et al., 2024) Overview of MU
(Li et al., 2025) Overview of MU
(Qu et al., 2024) Overview of LLM-U
(Liu et al., 2025b) Privacy and Security in MU
@@
Comparison with Existing Works. In recent years, numerous studies have surveyed the landscape of machine unlearning (MU), covering foundational frameworks (Nguyen et al., 2022a; Shaik et al., 2024; Xu, 2024; Li et al., 2024a; Zhang et al., 2023; Mercuri et al., 2022; Wang et al., 2024c; Jaman et al., 2024; Li et al., 2025), privacy and security risks (Liu et al., 2025a; Chen et al., 2025; Liu et al., 2025b), federated learning (Zhao et al., 2023; Romandini et al., 2024), generative models (Liu et al., 2024), and unlearning in large language models (Xu, 2024; Geng et al., 2025; Qu et al., 2024). However, as shown in Table 1, existing surveys generally lack a systematic and dedicated investigation into unlearning verification. In most of these works, verification is treated as an auxiliary component of model evaluation rather than as a standalone research focus. Although some studies touch on verification based on model performance (Zhao et al., 2023; Li et al., 2024a; Geng et al., 2025; Li et al., 2025), privacy analysis (Liu et al., 2025a; Zhang et al., 2023; Said et al., 2023; Chen et al., 2025; Wang et al., 2024c), and principles such as differential privacy or reproducibility (Zhao et al., 2023; Shaik et al., 2024; Mercuri et al., 2022), such discussions are often presented as background or peripheral remarks. They lack a unified framework for systematic categorization or cross-method comparison. More specifically, these surveys generally do not offer a structured comparison of different verification approaches across core dimensions such as theoretical foundations and evaluation accuracy. They also fail to examine practical considerations in real-world deployments, including feasibility, computational cost, and access constraints. Moreover, potential vulnerabilities and attack surfaces of existing verification techniques remain largely unexplored. These omissions hinder a comprehensive understanding of the capabilities, assumptions, and limitations of current unlearning verification mechanisms.
To bridge these gaps, this work presents the first systematic survey dedicated to verification methods in machine unlearning. We introduce a unified classification framework that encompasses both empirical and formal approaches, and assess existing methods across seven key dimensions: theoretical guarantees, access assumptions, sample-level granularity, verification accuracy, reliance on pre-injected data, efficiency and scalability, and method specificity. Our objective is to clarify the applicability and limitations of current verification techniques, and to provide a structured foundation for future research on verifiable machine unlearning in privacy-sensitive and compliance-driven settings.
@@ 2.1. Notation
$w\sim A(D)$ Original model
$w^{\prime}=M(w,D_{f})$ Unlearned model
@@
Figure 2. A framework diagram of machine unlearning
@@ 2.2. Machine Unlearning
Machine unlearning refers to the process of removing the influence of specific training samples from a previously trained machine learning model, while aiming to preserve the model’s original performance as much as possible, as shown in Figure 2. Formally, let $A$ be a randomized learning algorithm which, when trained on a dataset $D$, induces a distribution over model parameters. This defines a hypothesis space:
@@
Let $w\sim A(D)$ denote a specific model sampled from the distribution induced by training on $D$. We refer to $w$ as the original model, as it reflects the full influence of all training samples in $D$. Given a user request to eliminate the influence of a subset $D_{f}\subseteq D$, we refer to $D_{f}$ as the forgotten dataset.
To fulfill such a request, an unlearning mechanism $M$ is introduced. This mechanism operates on the original model $w$ and the forgotten dataset $D_{f}$, and produces a unlearned model:
$w^{\prime}=M(w,D_{f}),\quad w^{\prime}\in H^{\prime},$
@@
Based on the extent to which the influence of $D_{f}$ is removed, existing machine unlearning methods can be broadly classified into two categories: exact unlearning and approximate unlearning. Exact unlearning aims to completely forget $D_{f}$, requiring that the resulting model be indistinguishable from a model trained from scratch on the retained dataset $D\setminus D_{f}$, specifically in terms of its parameters. In comparison, approximate unlearning relaxes this requirement by allowing the resulting model to approximate the behavior or parameters of the retrained model. This relaxation enables improvements in computational efficiency, scalability, and real-world deployability.
@@ 2.2.2. Exact Unlearning
Let $A$ be a randomized learning algorithm, and let $M$ be an unlearning mechanism that transforms a specific model $w\sim A(D)$ into a new model $w^{\prime}=M(w,D_{f})$. Exact unlearning is satisfied if the distribution of the unlearned model $w^{\prime}$ matches that of retraining on $D_{r}$, that is,
$w^{\prime}\sim A(D_{r}).$
Equivalently, the unlearned model $w^{\prime}$ is required to lie within the hypothesis space
@@
which represents the set of models that can be obtained by training $A$ on the retained dataset. This condition ensures that the influence of $D_{f}$ is eliminated to the same extent as if those samples had never been included during training.
@@
SISA Methods. Bourtoule et al. (2021) propose the Sharded, Isolated, Sliced, and Aggregated (SISA) training framework, which partitions the training dataset into disjoint shards $D_{1},D_{2},\dots,D_{K}$, each trained independently to produce isolated sub-models $w^{1},w^{2},\dots,w^{K}$. The final model is then obtained by aggregating these sub-models:
$w=\texttt{Agg}(w^{1},w^{2},\dots,w^{K}).$
When a deletion request targets data $D_{f}\subseteq D_{j}$, only the affected shard $D_{j}$ is updated. The remaining data $D_{j}^{\prime}=D_{j}\setminus D_{f}$ is retrained to obtain an updated sub-model $w^{j^{\prime}}$, and the full model is re-aggregated as:
$w^{\prime}=\texttt{Agg}(w^{1},\dots,w^{j^{\prime}},\dots,w^{K}).$
@@
Subsequent works (Yan et al., 2022; Aldaghri et al., 2021; Kumar et al., 2022; Dukler et al., 2023; Golatkar et al., 2023; Chowdhury et al., 2024) extend the SISA design to improve unlearning efficiency and model performance. ARCANE (Yan et al., 2022) partitions the data by class and applies ensemble learning, transforming full-model retraining into per-class slice updates. Aldaghri et al. (Aldaghri et al., 2021) propose a coding-based sharding protocol, enabling lightweight sub-model recomputation through data-encoded representations. Modular adapter-based retraining (Kumar et al., 2022; Dukler et al., 2023; Golatkar et al., 2023) decouples update-sensitive components from the backbone model, while Chowdhury et al. (Chowdhury et al., 2024) introduces sequential slice training that enables localized fine-tuning. These designs aim to maintain the semantic guarantee of exact unlearning while minimizing the computational burden.
Non-SISA Methods. Non-SISA methods refer to machine unlearning approaches that do not adopt the ”Sharded, Isolated, Sliced, and Aggregated training” paradigm. Specifically, Cao et al. (Cao and Yang, 2015) propose a method that transforms the learning algorithm into a form dependent on a small number of summations, allowing training process to be based on these summations rather than individual data points. During the unlearning process, it only requires subtracting the influence of the deleted data from the summations and updating the model. Recently, Liu et al. (Liu et al., 2022b) introduce a fast retraining method using a distributed Quasi-Newton model update algorithm for data erasure, which is suitable for federated learning scenarios.

### v2

@@ Abstract.
With growing demands for privacy, security, and legal compliance (e.g., GDPR), machine unlearning has become a critical technique for ensuring the controllability of learning systems. A central challenge in this area is verifying whether unlearning has been successfully performed. Although unlearning methods are widely studied, verification remains underexplored and lacks a unified framework. This survey addresses the gap by organizing existing methods into behavioral and parametric categories based on the evidence used. It compares representative approaches in terms of assumptions, strengths, and vulnerabilities, and concludes with open problems to support the development of more reliable verification mechanisms.
@@
This work is supported by the National Natural Science Foundation of China (Grant No. 62372196).
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
As global concerns over data privacy continue to grow, particularly with the enforcement of regulations such as the General Data Protection Regulation (GDPR) (reg, 2018), managing data throughout its lifecycle has become a pressing compliance challenge. In response, machine unlearning has emerged as a key technology for protecting data sovereignty. By selectively removing specific data from trained models, it allows individuals to meaningfully exercise their right to be forgotten by enabling models to forget the influence of that data. For example, in facial recognition systems, when a user invokes their “right to be forgotten”, traditional forgotten methods merely remove stored records. In contrast, machine unlearning goes further by erasing the data’s influence from the model itself. Similarly, in intelligent customer service systems, when a user requests the deletion of chat history, the unlearning mechanism updates the model to forget the relevant interactions, ensuring that future responses are no longer shaped by the removed content.
@@
Existing machine unlearning techniques can be broadly classified into exact unlearning (Bourtoule et al., 2021; Yan et al., 2022; Kumar et al., 2022; Dukler et al., 2023; Golatkar et al., 2023; Chowdhury et al., 2025) and approximate unlearning (Felps et al., 2020; Chen et al., 2023; Tarun et al., 2023; Gou et al., 2021; Phuong and Lampert, 2019). Exact unlearning typically relies on retraining the model to fully eliminate the influence of specific data, which is computationally expensive. To improve the practicality of unlearning in real-world applications, a range of approximate unlearning methods have been proposed. These approaches strike a balance between effectiveness and efficiency, enabling models to approximate forgetting at a significantly lower computational cost.
@@
While algorithmic methods for machine unlearning offer mechanisms to erase specific training data from models, concerns about trust and transparency persist, particularly because most machine learning services operate as opaque, server-side systems. A central and unresolved question remains: how can model providers convincingly demonstrate that the targeted data has been thoroughly and irreversibly removed?
@@
In response, a diverse range of verification techniques has emerged in recent years, drawing from disciplines such as fingerprint (Zheng et al., 2022; Chai and Chen, 2022; Guo et al., 2021; An et al., 2024), privacy analysis (Tu et al., 2024; Di et al., 2025; Sharma et al., 2024), and performance-based diagnostics (Graves et al., 2021; Mehta et al., 2022; Chundawat et al., 2023a, b; Kurmanji et al., 2023). These approaches aim to evaluate whether a model retains residual information after an unlearning operation. However, the absence of a standardized verification framework has led to troubling inconsistencies. For instance, Jeon et al. (2024) demonstrates that a model may be deemed “successfully unlearned” under one verification criterion while simultaneously failing another—revealing a disconnect that undermines confidence in current practices.
@@
Overview. This paper presents a systematic overview of existing verification techniques for machine unlearning. The structure of this paper is illustrated in Figure 1. Specifically, Section 2 introduces the formal foundations of unlearning verification, including the definitions of exact and approximate unlearning, the construction of verification predicates, and the roles and interactions involved in the verification workflow. Section 3 reviews behavioral unlearning verification methods based on observable model behavior, whereas Section 4 covers parametric approaches that leverage internal model information and provide theoretical guarantees. Section 5 presents a comparative analysis of these approaches, examining their respective strengths and limitations. Section 6 investigates the reliability vulnerabilities of unlearning verification under adversarial settings, highlighting how weaknesses in parameter and behavioral signals may be exploited to deceive auditors. Finally, Section 7 outlines open problems in the field and highlights promising directions for future research.
@@
(Jaman et al., 2024) Overview of MU
(Li et al., 2024a) Overview of MU
(Li et al., 2025) Overview of MU
(Mercuri et al., 2022) Overview of MU
@@
(Shaik et al., 2024) Overview of MU
(Wang et al., 2024c) Overview of MU
(Zhang et al., 2023) Overview of MU
(Geng et al., 2025) Overview of LLM-U
(Qu et al., 2024) Overview of LLM-U
(Liu et al., 2024b) Overview of FU
(Romandini et al., 2024) Overview of FU
(Zhao et al., 2023) Overview of FU
(Xu, 2024) Overview of MU and LLM-U
(Said et al., 2023) Overview of GU
(Liu et al., 2024a) Overview of GenAI-U
(Chen et al., 2025) Privacy and Security in MU
(Liu et al., 2025b) Privacy and Security in MU
@@
Comparison with Existing Works. In recent years, numerous studies have surveyed the landscape of machine unlearning (MU), covering foundational frameworks (Nguyen et al., 2022a; Shaik et al., 2024; Xu, 2024; Li et al., 2024a; Zhang et al., 2023; Mercuri et al., 2022; Wang et al., 2024c; Jaman et al., 2024; Li et al., 2025), privacy and security risks (Liu et al., 2025a; Chen et al., 2025; Liu et al., 2025b), federated learning (Liu et al., 2024b; Zhao et al., 2023; Romandini et al., 2024), generative models (Liu et al., 2024a), and large language models (Xu, 2024; Geng et al., 2025; Qu et al., 2024). However, as shown in Table 1, existing surveys generally lack a systematic and dedicated investigation into unlearning verification. In most of these works, verification is treated as an auxiliary component of model evaluation rather than as a standalone research focus. Although some studies touch on verification based on model performance (Zhao et al., 2023; Li et al., 2024a; Geng et al., 2025; Li et al., 2025), privacy analysis (Liu et al., 2025a; Zhang et al., 2023; Said et al., 2023; Chen et al., 2025; Wang et al., 2024c), and principles such as differential privacy or reproducibility (Zhao et al., 2023; Shaik et al., 2024; Mercuri et al., 2022), such discussions are often presented as background or peripheral remarks. They lack a unified framework for systematic categorization or cross-method comparison. More specifically, these surveys generally do not offer a structured comparison of different verification approaches across core dimensions such as theoretical foundations and evaluation accuracy. They also fail to examine practical considerations in real-world deployments, including feasibility, computational cost, and access constraints. Moreover, potential vulnerabilities and attack surfaces of existing verification techniques remain largely unexplored. These omissions hinder a comprehensive understanding of the capabilities, assumptions, and limitations of current unlearning verification mechanisms.
To bridge these gaps, this work presents the first systematic survey dedicated to verification methods in machine unlearning. We introduce a unified classification framework that encompasses both behavioral and parametric approaches, and assess existing methods across seven key dimensions: theoretical guarantees, access assumptions, sample-level granularity, verification accuracy, reliance on pre-injected data, efficiency and scalability, and method specificity. Our objective is to clarify the applicability and limitations of current verification techniques, and to provide a structured foundation for future research on verifiable machine unlearning in privacy-sensitive and compliance-driven settings.
@@ 2.1. Notation
$\mathbf{w}^{*}$ The model that has never seen $D_{f}$
$\mathbf{w}$ Original model
$\mathbf{w}^{\prime}$ Unlearned model
@@
Figure 2. A framework diagram of machine unlearning.
@@ 2.2. Machine Unlearning
Machine unlearning refers to the process of removing the influence of specific training samples from a previously trained machine learning model, while preserving the model’s original performance as much as possible, as shown in Figure 2. Formally, let $A$ be a randomized learning algorithm which, when trained on a dataset $D$, induces a distribution over model parameters. This defines a hypothesis space:
@@
Let $\mathbf{w}\sim A(D)$ denote a specific model sampled from the distribution induced by training on $D$. We refer to $\mathbf{w}$ as the original model, as it reflects the full influence of all training samples in $D$. Suppose a user submits a request to eliminate the influence of a subset $D_{f}\subseteq D$, we refer to this subset $D_{f}$ as the forgotten dataset.
To fulfill such a request, an unlearning mechanism $M$ is introduced. This mechanism operates on the original model $\mathbf{w}$ and the forgotten dataset $D_{f}$, and produces a unlearned model:
$\mathbf{w}^{\prime}=M(\mathbf{w},D_{f}),\quad\mathbf{w}^{\prime}\in H^{\prime},$
@@
Based on the extent to which the influence of $D_{f}$ is removed, existing machine unlearning methods can be broadly classified into two categories: exact unlearning and approximate unlearning. Exact unlearning aims to completely forget $D_{f}$, requiring that the resulting model be indistinguishable from a model trained from scratch on the retained dataset $D\setminus D_{f}$, specifically in terms of its parameters. In comparison, approximate unlearning relaxes this requirement by allowing the resulting model to approximate the behavior or parameters of the retrained model. This relaxation enables improvements in computational efficiency, scalability, and real-world deployability. Table 3 summarizes representative methods across both categories, which will be discussed in detail below.
@@ 2.2.2. Exact Unlearning
Let $A$ be a randomized learning algorithm, and let $M$ be an unlearning mechanism that transforms a specific model $\mathbf{w}\sim A(D)$ into a new model $\mathbf{w}^{\prime}=M(w,D_{f})$. Exact unlearning is satisfied if the distribution of the unlearned model $\mathbf{w}^{\prime}$ matches that of retraining on $D_{r}$, that is,
$\mathbf{w}^{\prime}\sim A(D_{r}).$
Equivalently, the unlearned model $\mathbf{w}^{\prime}$ is required to lie within the hypothesis space
@@
which comprises all models attainable when $A$ is trained solely on the retained data. This condition ensures that the influence of $D_{f}$ is eliminated to the same extent as if those samples had never been included during training.
@@
SISA Methods. Bourtoule et al. (2021) proposed the Sharded, Isolated, Sliced, and Aggregated (SISA) training framework, which partitions the training dataset into disjoint shards $D_{1},D_{2},\dots,D_{K}$, each trained independently to produce isolated sub-models $\mathbf{w}^{1},\mathbf{w}^{2},\dots,\mathbf{w}^{K}$. The final model is then obtained by aggregating these sub-models:
$\mathbf{w}=\texttt{Agg}(\mathbf{w}^{1},\mathbf{w}^{2},\dots,\mathbf{w}^{K}).$
When a deletion request is issued for a subset $D_{f}\subseteq D_{j}$, only the affected shard $D_{j}$ is updated. The remaining data $D_{j}^{\prime}=D_{j}\setminus D_{f}$ is retrained to obtain an updated sub-model $\mathbf{w}^{j^{\prime}}$, and the full model is re-aggregated as:
$\mathbf{w}^{\prime}=\texttt{Agg}(\mathbf{w}^{1},\dots,\mathbf{w}^{j^{\prime}},\dots,\mathbf{w}^{K}).$
@@
Subsequent works (Yan et al., 2022; Aldaghri et al., 2021; Kumar et al., 2022; Dukler et al., 2023; Golatkar et al., 2023; Chowdhury et al., 2025) extend the SISA design to improve unlearning efficiency and model performance. ARCANE (Yan et al., 2022) partitions the data by class and applies ensemble learning, transforming full-model retraining into per-class slice updates. Aldaghri et al. (Aldaghri et al., 2021) proposed a coding-based sharding protocol, enabling lightweight sub-model recomputation through data-encoded representations. Modular adapter-based retraining (Kumar et al., 2022; Dukler et al., 2023; Golatkar et al., 2023) decouples update-sensitive components from the backbone model, while Chowdhury et al. (Chowdhury et al., 2025) introduced sequential slice training that enables localized fine-tuning. These designs aim to maintain the semantic guarantee of exact unlearning while minimizing the computational burden.
Non-SISA Methods. Non-SISA methods refer to machine unlearning approaches that do not adopt the ”Sharded, Isolated, Sliced, and Aggregated” training paradigm. Specifically, Cao et al. (Cao and Yang, 2015) proposed a method that transforms the learning algorithm into a form dependent on a small number of summations, allowing training process to be based on these summations rather than individual data points. During the unlearning process, it only requires subtracting the influence of the deleted data from the summations and updating the model. Recently, Liu et al. (Liu et al., 2022b) introduced a fast retraining method using a distributed Quasi-Newton model update algorithm for data erasure, which is suitable for federated learning scenarios.